i) A particle moves along a line with velocity function v(t)= (t2 - t), where v is measured in meters per second. Find (a) the displacement and (b) the distance traveled by

the particle during the time interval [ 0, 5]


ii) Use the properties of integrals to verify the inequality.

Integral of "\intop (sinx/x) dx" Upper limit ("\pi" /2) and Lower limit ("\pi" /4) <= ( √2 / 2) .


iii).Evaluate the intregal, if it exists.

  1. "\intop (1+ tan t)" 3 sec2t dt ( Upper limit ("\pi" /4) and Lower limit (0)
  2. "\intop\begin{vmatrix}
 √x - 1 \\

\end{vmatrix}" dx (Upper limit (4) and Lower limit (0)
  3. "\intop"tan x ln(cos x) dx
1
Expert's answer
2021-09-07T17:29:04-0400

i)

a)


"s(t)=\int v(t)dt=\int(t^2-t)dt=\dfrac{t^3}{3}-\dfrac{t^2}{2}+s(0),m"



b)


"s(5)=\dfrac{(5)^3}{3}-\dfrac{(5)^2}{2}+s(0)=\dfrac{175}{6}+s(0) (m)"

"distance=s(5)-s(0)=\dfrac{175}{6}+s(0) -s(0)"

"=\dfrac{175}{6}(m)"

ii)


"\sin x=\displaystyle\sum_{i=0}^{\infin}\dfrac{(-1)^{n}x^{2n+1}}{(2n+1)!}"

"\dfrac{\sin x}{x}=\displaystyle\sum_{i=0}^{\infin}\dfrac{(-1)^{n}x^{2n}}{(2n+1)!}"

"=1-\dfrac{x^2}{3!}+\dfrac{x^4}{5!}-\dfrac{x^6}{7!}+..."

Integrate term by term


"\int\dfrac{\sin x}{x}dx=\int\displaystyle\sum_{i=0}^{\infin}\dfrac{(-1)^{n}x^{2n}}{(2n+1)!}dx"

"=\displaystyle\sum_{i=0}^{\infin}\dfrac{(-1)^{n}x^{2n+1}}{(2n+1)!(2n+1)}+C"

"=C+x-\dfrac{x^3}{3!(3)}+\dfrac{x^5}{5!(5)}-\dfrac{x^7}{7!(7)}+..."

Use the second fundamental theorem of calculus to evaluate


"\displaystyle\int_{\pi/4}^{\pi/2}\dfrac{\sin x}{x}dx=\displaystyle\int_{\pi/4}^{\pi/2}\displaystyle\sum_{i=0}^{\infin}\dfrac{(-1)^{n}x^{2n}}{(2n+1)!}dx"

"=\bigg[x-\dfrac{x^3}{3!(3)}+\dfrac{x^5}{5!(5)}-\dfrac{x^7}{7!(7)}+...\bigg]\begin{matrix}
 \pi/2 \\
 \pi/4
\end{matrix}"

"\leq\dfrac{\pi}{2}-\dfrac{(\dfrac{\pi}{2})^3}{18}+\dfrac{(\dfrac{\pi}{2})^5}{600}-(\dfrac{\pi}{4}-\dfrac{(\dfrac{\pi}{4})^3}{18}+\dfrac{(\dfrac{\pi}{4})^5}{600})"

"\leq0.6125\leq\dfrac{\sqrt{2}}{2}"

Therefore

"\displaystyle\int_{\pi/4}^{\pi/2}\dfrac{\sin x}{x}dx\leq\dfrac{\sqrt{2}}{2}"

iii)

1.


"\int(1+\tan t)^3\sec^2tdt"

"u=1+\tan t, du=\sec^2t dt"

"\int(1+\tan t)^3\sec^2tdt=\int u^3du=\dfrac{u^4}{4}+C"

"=\dfrac{(1+\tan t)^4}{4}+C"

2.


"\displaystyle\int_{0}^{4}|\sqrt{x}-1|dx=-\displaystyle\int_{0}^{1}(\sqrt{x}-1)dx+\displaystyle\int_{1}^{4}(\sqrt{x}-1)dx"

"=-[\dfrac{2x^{3/2}}{3}-x]\begin{matrix}
 1 \\
 0
\end{matrix}+[\dfrac{2x^{3/2}}{3}-x]\begin{matrix}
 4 \\
 1
\end{matrix}"

"=-\dfrac{2}{3}+1+0+\dfrac{16}{3}-4-\dfrac{2}{3}+1=2"

3.


"\int \tan x\ln(\cos x)dx"

"u=\ln(\cos x), du=\dfrac{1}{\cos x}(-\sin x)dx=-\tan xdx"

"\int \tan x\ln(\cos x)dx=-\int udu=-\dfrac{u^2}{2}+C"

"=-\dfrac{\ln^2(\cos x)}{2}+C"